OpenStudy (anonymous):

please show me how to factor completely 3y^2-42y-96

OpenStudy (anonymous):

factor a 3 out 3(y^2-14y-32) try solving it. if you have trouble just ask and I can help you with the next step.

OpenStudy (anonymous):

im so confused i dont even know where to start i know the common is 4

OpenStudy (anonymous):

ok i was so wrong common is not 4

OpenStudy (anonymous):

3(y-16)(y+2) -16+2 = 14 -16*2 = -32

OpenStudy (anonymous):

write the middle term as the sum of 2 terms in such a way that you will have a common factor.

OpenStudy (anonymous):

im gonna try another one and see if i can get it

OpenStudy (anonymous):

that's a good idea.

OpenStudy (anonymous):

10-19s+6s would this one be considered prime there are no numbers that go into 19?